# Non-Markovian dynamics in nonstationary Gaussian baths

### v2

@@ Abstract
Building on the standard hierarchy of pure states (HOPS) approach, we construct a generalized formulation suitable for open quantum systems interacting with nonstationary Gaussian baths, potentially extending its applicability to nonequilibrium baths. This is achieved by extending the conventional exponential decomposition of bath correlation functions (BCF) to allow explicitly time-dependent forms. We demonstrate the method’s performance on two examples of nonstationary squeezed reservoirs generated via uniform squeezing and degenerate parametric amplification. Benchmarking against the associated hierarchy of master equations shows that HOPS achieves superior efficiency under hierarchy truncation. In cases where each contribution in the BCF expansion can be associated with an independent physical bath, the formalism can be simplified in a pseudomode representation which is more efficient in a strongly non-Markovian regime. Our results highlight HOPS as a versatile and powerful tool for simulating open quantum systems in nonstationary baths, with potential applications ranging from squeezed light-matter interactions to driven quantum materials and dissipative phase transitions.
@@ I Introduction
Most such methods evolve density operators. However, it is often more efficient to sample stochastic pure-state trajectories and reconstruct the density matrix via ensemble averaging [25, 37]. A notable method in this direction is the hierarchy of pure states (HOPS) approach [38, 39, 40, 41], which assumes exponential decomposition of the BCF and unravels non-Markovian dynamics into a set of stochastic differential equations for auxiliary states. This method effectively halves the number of degrees of freedom and, in some cases, has been shown to converge faster with respect to the number of hierarchy levels than approaches based on hierarchies of density operators [42].
@@ II Theory
$\hat{H}(t)=\hat{H}_{S}+\hbar\hat{L}\hat{B}(t),$ (1)
where $\hat{H}_{S}$ is the Hamiltonian of the isolated system, which may be time-dependent. The second term describes system-bath coupling: $\hat{L}$ and $\hat{B}(t)$ are Hermitian operators of the system and the bath, respectively.
@@ II.1 Hierarchy of pure states (HOPS)
When the bath is stationary and characterized by a time-translation invariant BCF, $\alpha(t,s)=\alpha(t-s)$, the reduced system dynamics can be efficiently treated using the hierarchy of pure states (HOPS) [38]. In this method, the continuous bath spectrum is approximated by a finite set of effective broadband modes. The reduced system dynamics is then described in terms of stochastic pure states, reducing computational cost.
@@
Extensions of HOPS to nonstationary BCFs have recently been explored [43]. In this article, we extend HOPS to a broad class of time-dependent baths with BCFs that depend on both time arguments, $\alpha(t,s)\neq\alpha(t-s)$. Specifically, we generalize the exponential ansatz to
@@
The functions $f_{j}(t)$ and $g_{j}(s)$ introduce explicit time dependence, thereby breaking time-translation invariance and enabling the treatment of such baths.
In the following paragraphs, we present the resulting equations for HOPS. Detailed derivations are provided in Appendix A. The system’s reduced density matrix $\hat{\rho}_{S}(t)$ is obtained as a statistical average over stochastic pure states $|\psi(t)\rangle$:
@@
where $\hat{O}$ denotes an operator acting on the system. This formulation replaces the reduced density matrix with pure states, thereby reducing the number of degrees of freedom by half.
@@
In this representation, truncating the hierarchy is equivalent to truncating the pseudo-Fock basis. Auxiliary states with $\mathbf{n}\neq\mathbf{0}$ encode bath properties [47].
The extended state vector $|\Psi(t)\rangle$ is governed by a stochastic Schrödinger equation
@@
The second term in Eq. (9) contains the effective Hamiltonian $\hat{H}_{\text{eff}}(t)$, which includes the isolated system Hamiltonian $\hat{H}_{S}$ and interaction with the effective modes:
$\hat{H}_{\text{eff}}(t)=\hat{H}_{S}+\hbar\sum_{j=1}^{N}\sqrt{\dfrac{\Gamma_{j}}{2}}\big\{f_{j}(t)\hat{c}_{j}+g^{*}_{j}(t)\hat{c}_{j}^{\dagger}\big\}\hat{L}$. (10)
@@ II.2 Nonlinear stochastic Schrödinger equation
Monte Carlo evaluation of observables using Eq. (7) can converge slowly when using solutions of the linear equation (9), especially in the strong coupling regime [39]. This issue can be mitigated by applying the Girsanov transformation [38, 39, 25], which normalizes each stochastic realization when computing observables, thus ensuring equal weight for all realizations and improving sampling efficiency.
After applying the Girsanov transformation, the expression for the reduced density matrix takes the form
@@
The nonlinearity in Eq. (13) originates parly from the last term, which involves the normalized expectation value of $\hat{L}$ for a given realization of $|\tilde{\psi}(t)\rangle$:
@@
which includes a deterministic shift. The integral can be efficiently computed by solving a system of differential equations. In the special case where $f_{j}(t)=g_{j}(t)$, the entire $\tilde{Z}(t)$ can be obtained by solving a set of stochastic differential equations.
@@ II.3 Hierarchy of master equations (HME)
This deterministic treatment avoids stochastic sampling but doubles the number of degrees of freedom. The density operator $\hat{\rho}(t)$ evolves according to the following master equation: [^1] This derivation uses the relation $\mathbb{E}\big[Z^{*}(t)|\Psi(t)\rangle\langle\Psi(t)|\big]=\displaystyle\sum_{j=1}^{N}\sqrt{\frac{\Gamma_{j}}{2}}f_{j}(t)\hat{c}_{j}\hat{\rho}(t)$ and its complex conjugate.
@@ II.4 Pseudomode master equation (PME)
The computational complexity of solving Eq. (21) can be reduced by half, at the cost of introducing statistical sampling through the standard Markovian quantum-state diffusion formalism [25, 52, 53]. The corresponding white-noise pseudomode stochastic Schrödinger equation (PSSE) is given in Appendix C, and will be used in our numerical analysis.
@@ III.1 Single mode BCF
The atomic dynamics induced by the bath with the correlation function in Eq. (22) can also be treated using the pseudomode approach described in Sec. II.4. Thus, this model provides an ideal test case for benchmarking and comparing the performance of all methods considered in this article. For the analysis, we fix the coupling strength $\gamma=1$, central frequency $\omega_{0}=5$, and squeezing parameters $r=1.5$ and $\varphi=0$, while varying the spectral half-width $\Gamma$. Figure 1 shows the time evolution of the mean Bloch vector components for $\Gamma=0.2$ (a), $0.5$ (b), $1.0$ (c), and $2.0$ (d). Increasing $\Gamma$ leads to a faster decay of the transverse Bloch vector components, while the $z$ component exhibits oscillations due to driving. The bottom row (e)-(h) displays the corresponding components in the rotating frame, highlighting that the $x$ and $y$ components decay at different rates. In this plot, we numerically solve the pseudomode master equation (21) with $n^{\text{max}}=100$. In practice, significantly fewer hierarchies are often sufficient to reach the desired accuracy.
@@
Figure 2 presents the estimated numerical error, which includes contributions from hierarchy truncation and discretization due to the finite time step [see Appendix D for error expressions]. The figure compares the numerical errors of HME (18) and PME (21) in the truncated pseudo-Fock space. The top row shows the time dependence of the error for the HME. As can be seen in panel (a), simulating longer times may require a larger number of hierarchies. In panels (b)-(d), as the parameter $\Gamma$ increases, the time after which the error remains effectively constant occurs earlier, provided that $n^{\text{max}}$ is sufficiently large. The time dependence of the error for the PME exhibits qualitatively similar behavior and is not shown. The bottom row presents the time-averaged root-mean-square error for both methods. In panel (e), for $\Gamma=0.2$, the PME achieves better accuracy with a much smaller truncation level. For $\Gamma=0.5$ (f), the error for both methods becomes comparable after $30$ hierarchies. In the near-Markovian regime shown in panels (g) and (h), the HME outperforms the PME. Increasing the number of hierarchies in the HME may increase the error, as clearly visible in panels (e)–(g). In contrast, the PME remains stable, since including more hierarchies does not increase the error. In all cases, the error curves eventually saturate, approaching a constant baseline determined by the fourth-order Runge-Kutta accuracy.
The corresponding analysis for the Girsanov‐transformed HOPS and PSSE is shown in Fig. 3. Expressions for error estimation are provided in Appendix E. The top row illustrates the time dependence of the error for HOPS. As in Fig. 2, increasing $\Gamma$ shifts the constant‐error plateau to earlier times. The PSSE displays qualitatively similar behavior and is not shown. The bottom row shows root-mean-square error. For $\Gamma=0.2$ in panel (e), the PSSE achieves higher accuracy with a much smaller truncation level. At $\Gamma=0.5$ in panel (f), the error of both methods becomes comparable after moderate truncation levels. In the near‐Markovian regime in panels (g)-(h), the hierarchy of pure states outperforms the pseudomode representation. Unlike the HME, which can become unstable for small truncation levels, both stochastic approaches remain stable, with the error either decreasing or saturating at the baseline defined by Monte Carlo sampling. Panels (e)-(h) also include the mean master-equation errors from Fig. 2 for comparison. While the PSSE error closely follows the PME error until it reaches the Monte Carlo baseline, the HOPS error differs significantly from the HME error. This difference arises because the ensemble average of the extended state vector in a truncated pseudo-Fock space does not yield the corresponding truncated density operator,
$\mathbb{E}\big[|\Psi_{\text{trunc}}(t)\rangle\langle\Psi_{\text{trunc}}(t)|\big]\not=\hat{\rho}_{\text{trunc}}(t),$
as the derivation of Eq. (18) relies on retaining the full, untruncated basis. For the BCF considered here, HOPS is more stable and less sensitive to the truncation level, since certain bath features are also encoded in the noise term $Z(t)$.
@@ III.2 Three mode BCF
which is composed of three effective modes with decay rates $\Gamma_{1}=\Gamma_{0}$, $\Gamma_{2}=\Gamma-\epsilon$, and $\Gamma_{3}=\Gamma+\epsilon$. Here, $\alpha_{j}(\tau)$ are exponential kernels [see Eq. (4)]. The nonstationary functions $f_{i}(t)$ are
@@ A.1 Tracing out the bath in a coherent-state representation
$\dfrac{\partial|\psi(\bm{z}^{*},t)\rangle}{\partial t}=\Big[-\dfrac{i}{\hbar}\hat{H}_{S}-i\mathcal{D}(t)\hat{L}\\ -iZ^{*}(t)\hat{L}\Big]|\psi(\bm{z}^{*},t)\rangle.$ (40)
@@ A.2 Eliminating the functional derivative: a hierarchy
Note that the derivative of the full operator $\mathcal{D}(t)$ cannot be expressed solely in terms of $\mathcal{D}(t)$. Thus, constructing the hierarchy from $\mathcal{D}(t)$ does not yield a closed set of equations, and instead we use the components $\mathcal{D}_{j}(t)$.
@@ Appendix C Pseudomode stochastic Schrödinger equation (PSSE)
where $S_{i}(t)$ are complex Gaussian white noises with zero mean and correlations:
@@
Note that when the rates $\Gamma_{j}$ are close to zero, the overall state is close to a pure state, which makes the statistical sampling less relevant leading to better convergence.
@@
$\tilde{S}_{j}(t)=S_{j}(t)+\sqrt{2\Gamma_{j}}\dfrac{\langle\tilde{\Psi}^{\prime}(t)|\hat{c}^{\dagger}_{i}|\tilde{\Psi}^{\prime}(t)\rangle}{\langle\tilde{\Psi}^{\prime}(t)|\tilde{\Psi}^{\prime}(t)\rangle}.$ (62)

### v3

@@ Abstract
Building on the standard hierarchy of pure states (HOPS) approach, we construct a generalized formulation suitable for open quantum systems interacting with nonstationary Gaussian baths, potentially extending its applicability to nonequilibrium baths. This is achieved by extending the conventional exponential decomposition of a bath correlation function (BCF) for nonstationary cases. Using our formulation of HOPS, we derive the corresponding hierarchy of master equations and, when each term in the BCF expansion can be associated with an independent physical bath, we show how the formalism connects to the well-known pseudomode representation. We demonstrate the method’s performance on two examples of nonstationary squeezed reservoirs generated via uniform squeezing and degenerate parametric amplification in a one-sided cavity. Benchmarking against the hierarchy of master equations shows that HOPS is more efficient under hierarchy truncation. The pseudomode representation is shown to be more efficient in the strongly non-Markovian regime. Our results highlight HOPS as a versatile and powerful tool for simulating open quantum systems in nonstationary baths, with potential applications ranging from squeezed light-matter interactions to driven quantum materials and dissipative phase transitions.
@@ I Introduction
Most such methods evolve density operators. However, it is often more efficient to sample stochastic pure-state trajectories and reconstruct the density matrix via ensemble averaging [18, 62]. A notable method in this direction is the hierarchy of pure states (HOPS) approach [64, 34, 35, 30], which assumes exponential decomposition of the BCF and unravels non-Markovian dynamics into a set of stochastic differential equations for auxiliary states. This method reduces the dimensionality of the state space from Liouville space to Hilbert space and, in some cases, has been shown to converge faster with respect to the number of hierarchy levels than approaches based on hierarchies of density operators [74].
@@ II Theory
$\hat{H}(t)=\hat{H}_{S}(t)+\hbar\hat{L}\hat{B}(t),$ (1)
where $\hat{H}_{S}(t)$ is the Hamiltonian of the isolated system, which may be time-dependent. The second term describes system-bath coupling: $\hat{L}$ and $\hat{B}(t)$ are Hermitian operators of the system and the bath, respectively.
@@ II.1 Hierarchy of pure states (HOPS)
When the bath is stationary and characterized by a time-translation invariant BCF, $\alpha(t,s)=\alpha(t-s)$, the reduced system dynamics can be efficiently treated using the hierarchy of pure states (HOPS) [64]. In this method, the continuous bath spectrum is approximated by a finite set of effective broadband modes. The reduced system dynamics is described by averaging over stochastic states, rather than propagating density operators of larger dimension.
@@
Extension of HOPS to nonstationary BCFs has recently been explored in Ref. [42], which considered a specific BCF for a squeezed reservoir. In this article, we consider a more general class of nonstationary BCFs, $\alpha(t,s)\neq\alpha(t-s)$, of the form
@@
The functions $f_{j}(t)$ and $g_{j}(s)$ introduce explicit time dependence. They cannot be chosen independently; there is an implicit relation between them due to the Hermiticity and positive semidefiniteness of the BCF.
In Appendix A, we provide our derivation of the hierarchy of pure states, adapted for the BCF in Eq. (5). We use the pseudo-Fock basis [24] to obtain a compact form of the equations of motion, and present only the final equations below.
The system’s reduced density matrix $\hat{\rho}_{S}(t)$ is obtained as a statistical average over stochastic pure states $|\psi(t)\rangle$:
@@
where $\hat{O}$ denotes an operator acting on the system. Reformulating the problem in terms of stochastic pure states reduces the dimensionality of the state space. This partially mitigates the bottleneck associated with the unfavorable scaling of Liouville space by trading it for statistical sampling, which can be easily parallelized. The main practical limitation instead lies in the convergence of the statistical averages, which is addressed later.
@@
Truncating the hierarchy is equivalent to truncating the pseudo-Fock basis. The physical state is normalized on average, whereas the norms of the auxiliary states encode normal-ordered expectation values of the bath operator [4].
The key theoretical result of our work is a stochastic Schrödinger equation for the extended state vector
@@
The second term in Eq. (9) contains the effective Hamiltonian $\hat{H}_{\text{eff}}(t)$, which includes the isolated system Hamiltonian $\hat{H}_{S}(t)$ and interaction with the effective modes:
$\hat{H}_{\text{eff}}(t)=\hat{H}_{S}(t)+\hbar\sum_{j=1}^{N}\sqrt{\dfrac{\Gamma_{j}}{2}}\big\{f_{j}(t)\hat{c}_{j}+g^{*}_{j}(t)\hat{c}_{j}^{\dagger}\big\}\hat{L}$. (10)
@@ II.2 Nonlinear stochastic Schrödinger equation
Monte Carlo evaluation of observables using Eq. (7) can converge slowly when using the solutions of the linear equation (9), especially in the strong-coupling regime [34]. This can be improved with a Girsanov transformation [64, 34, 18]. First, the stochastic state’s fluctuating norm is absorbed into the sampling measure, so observables are computed from normalized quantum trajectories. Second, the sampling is guided toward physically relevant regions where the state is localized [18].
The reduced density matrix then takes the form
@@
The nonlinearity in Eq. (13) originates partly from the last term, which involves the normalized expectation value of $\hat{L}$ for a given realization of $|\tilde{\psi}(t)\rangle$:
@@
which includes an additional term that can be viewed as an analogue of the conditional polarization field generated by the system for a given stochastic realization. Consequently, the noise Z(t){Z}(t) appearing in the first term—whose correlation properties remain unchanged—represents fluctuations about this conditional field. The integral can be efficiently computed by solving a system of differential equations. In the special case where $f_{j}(t)=g_{j}(t)$, the entire $\tilde{Z}(t)$ can be obtained by solving a set of stochastic differential equations.
@@ II.3 Hierarchy of master equations (HME)
This deterministic treatment avoids stochastic sampling but increases the size of the state space. The density operator $\hat{\rho}(t)$ evolves according to the following master equation: [^1] This derivation uses the relation $\mathbb{E}\big[Z(t)|\Psi(t)\rangle\langle\Psi(t)|\big]=\displaystyle\sum_{j=1}^{N}\sqrt{\frac{\Gamma_{j}}{2}}f_{j}(t)\hat{c}_{j}\hat{\rho}(t)$ and its Hermitian conjugate.
@@
As a side remark, the master equation (18) can, in principle, be unraveled using stochastic states driven by white noise, provided one enlarges the state space by allowing the stochastic bra and ket to differ and obey different equations. Related approaches have been considered in Refs. [47, 39, 5]. However, for the problems considered in Sec. III, we found that the resulting equations converge poorly. In addition, when we normalized individual trajectories, some realizations became unstable and exhibited unbounded growth, and we were unable to stabilize the dynamics. For these reasons, we do not pursue this approach further here and provide no equations or figures. This behavior is reminiscent of the runaway trajectories encountered within the positive-$P$ representation [31, 17, 10, 9, 11]. This comparison highlights an advantage of the colored-noise unraveling provided by HOPS in Eq.  (13): it evolves a single stochastic vector (the bra and ket are conjugate) and remains stable.
@@ II.4 Pseudomode master equation (PME)
The computational complexity of solving Eq. (21) in Liouville space can be reduced to Hilbert-space scaling, at the cost of introducing statistical sampling through the standard Markovian quantum-state diffusion formalism [18, 32, 29]. The corresponding white-noise pseudomode stochastic Schrödinger equation (PSSE) is given in Appendix C, and will be used in our numerical analysis.
@@ III.1 Single mode BCF
The atomic dynamics induced by the bath with the correlation function in Eq. (22) can also be treated using the pseudomode approach described in Sec. II.4. Thus, this model provides an ideal test case for benchmarking and comparing the performance of all methods considered in this article. For the analysis, we fix the coupling strength $\gamma=1$, central frequency $\omega_{0}=5$, and squeezing parameters $r=1.5$ and $\varphi=0$, while varying the spectral half-width $\Gamma$.
Figure 1 shows the time evolution of the mean Bloch vector components for $\Gamma=0.2$ (a), $0.5$ (b), $1.0$ (c), and $2.0$ (d). The atom evolves into a nonstationary state with oscillations in the mean $z$ component. Increasing $\Gamma$ leads to a faster decay of the transverse Bloch vector components. The bottom row (e)-(h) displays the corresponding components in the rotating frame, highlighting that the $x$ and $y$ components decay at different rates—an effect of squeezing [26]. In this plot, we numerically solve the pseudomode master equation (21) with $n^{\text{max}}=100$. In practice, significantly fewer hierarchies are often sufficient to reach the desired accuracy. In what follows, we show that a finite number of hierarchies is sufficient to reproduce this long-time nonstationary dynamics.
Figure 2 compares the numerical errors of the HME (18) and the PME (21) solved in the truncated pseudo-Fock space. The estimated error includes both the hierarchy truncation error and discretization error arising from the finite time step [see Appendix D for error expressions].
The top row (a)-(d) shows the time dependence of the error for the HME. As can be seen in panel (a), simulating longer times may require a larger number of hierarchies. In panels (b)-(d), as the parameter $\Gamma$ increases, the time after which the error remains effectively constant occurs earlier, provided that $n^{\text{max}}$ is sufficiently large. The time dependence of the error for the PME exhibits similar behavior and is not shown.
@@
The bottom row (e)-(h) presents the time-averaged root-mean-square error for both methods. In panel (e), for $\Gamma=0.2$, the PME achieves better accuracy with a much smaller truncation level. For $\Gamma=0.5$ (f), the error for both methods becomes comparable after $30$ hierarchies. In the near-Markovian regime shown in panels (g) and (h), the HME requires fewer hierarchies than the PME. As we move closer to the non-Markovian regime in panel (e), the HME error does not decrease monotonically with increasing the number of hierarchies: it first grows and only then decreases. By contrast, the PME remains stable—adding more hierarchy levels does not increase the error. In all cases, the error curves eventually saturate, approaching a constant baseline determined by the fourth-order Runge-Kutta accuracy.
The corresponding analysis for the Girsanov‐transformed HOPS and PSSE is shown in Fig. 3. Expressions for error estimation are provided in Appendix E. The top row (a)-(d) illustrates the time dependence of the error for HOPS: increasing $\Gamma$ shifts the constant‐error plateau to earlier times. The PSSE displays qualitatively similar behavior and is not shown.
The bottom row (e)-(h) shows root-mean-square error. For $\Gamma=0.2$ in panel (e), the PSSE achieves higher accuracy with a much smaller number of hierarchies. At $\Gamma=0.5$ in panel (f), the error of both methods becomes comparable after moderate truncation levels. In the near‐Markovian regime in panels (g)-(h), the HOPS outperforms the PSSE.
Unlike the HME, which demonstrated instability at small truncation levels, both stochastic approaches remain stable, with the error either decreasing or saturating at the baseline defined by Monte Carlo sampling. Panels (e)-(h) also include the mean master-equation errors from Fig. 2 for comparison. The PSSE error closely follows the PME error until it reaches the Monte Carlo baseline. This behavior is expected, as ensemble average of stochastic state vectors in the truncated space converges to truncated solution of the PME.
In contrast, the HOPS error differs substantially from the HME error. This mismatch can be expected because, in a truncated pseudo-Fock space, the ensemble average of HOPS stochastic projectors generally does not reproduce the truncated HME density operator, $\mathbb{E}\big[|\Psi_{\text{trunc}}(t)\rangle\langle\Psi_{\text{trunc}}(t)|\big]\not=\hat{\rho}_{\text{trunc}}(t)$, since the derivation of the HME (18) relies on retaining the full, untruncated basis. The stability of HOPS—and its higher tolerance to hierarchy truncation compared with the HME—can be attributed to the fact that part of the bath influence is carried by the stochastic term $Z(t)$, in addition to the hierarchy itself. As a result, truncating the pseudo-Fock space has a weaker impact, because the statistical properties of $Z(t)$ do not depend on the truncation. We emphasize that our HME benchmark was performed only for the specific bath-correlation function considered here; therefore, we cannot draw general conclusions about the stability of the HME.
@@ III.2 Three mode BCF
which is composed of three effective modes with decay rates $\Gamma_{1}=\Gamma_{0}$, $\Gamma_{2}=\Gamma-\epsilon$, and $\Gamma_{3}=\Gamma+\epsilon$. Here, $\alpha_{j}(\tau)$ are exponential kernels [see Eq. (4)]. The nonstationary functions $f_{i}(t)$ are given by
@@ A.1 Tracing out the bath in a coherent-state representation
$\dfrac{\partial|\psi(\bm{z}^{*},t)\rangle}{\partial t}=\Big[-\dfrac{i}{\hbar}\hat{H}_{S}(t)-i\mathcal{D}(t)\hat{L}\\ -iZ^{*}(t)\hat{L}\Big]|\psi(\bm{z}^{*},t)\rangle.$ (40)
@@ A.2 Eliminating the functional derivative: a hierarchy
Note that the derivative of the full operator $\mathcal{D}(t)$ cannot be expressed solely in terms of $\mathcal{D}(t)$. Thus, constructing the hierarchy from $\mathcal{D}(t)$ does not yield a closed set of equations, thus, we use instead the components $\mathcal{D}_{j}(t)$.
@@ Appendix C Pseudomode stochastic Schrödinger equation (PSSE)
where $S_{i}(t)$ are complex Gaussian white noise terms with zero mean and correlations:
@@
Note that when the rates $\Gamma_{j}$ are close to zero, the overall state is close to a pure state, which makes the statistical sampling less relevant, thereby leading to better convergence.
@@
$\tilde{S}^{*}_{j}(t)=S^{*}_{j}(t)+\sqrt{2\Gamma_{j}}\dfrac{\langle\tilde{\Psi}^{\prime}(t)|\hat{c}^{\dagger}_{i}|\tilde{\Psi}^{\prime}(t)\rangle}{\langle\tilde{\Psi}^{\prime}(t)|\tilde{\Psi}^{\prime}(t)\rangle}.$ (62)